# Temporally-Consistent Koopman Autoencoders for Forecasting Dynamical Systems

### v2

@@ Methods
The experiments were carried out on the Ohio Supercomputer’s (OSC) [42] Owens cluster, consisting of 842 Dell Nodes and 648 compute nodes with Dell PowerEdge C6320 two-socket servers, Intel Xeon E5-2680 v4 (Broadwell, 14 cores, 2.40 GHz) processors and 128 GB memory. As shown in Figure 1, we partition the dataset $\mathbf{X}$ into training ($\mathbf{X}_{\text{train}}$), validation ($\mathbf{X}_{\text{val}_{1}},\mathbf{X}_{\text{val}_{2}}$) and testing ($\mathbf{X}_{\text{test}}$) set. For a set of hyperparameters, the tcKAE architecture is trained to reduce the training loss $\mathcal{L}_{\text{tot}}$ as defined in 12. The set of hyperparameters which provides lowest validation loss is chosen to be the optimal set of hyperparameters for each case. The validation loss is computed as the average relative 2-norm error over $\mathbf{X}_{\text{val}_{2}}$, using predictions initialized from points in $\mathbf{X}_{\text{val}_{1}}$ as initial conditions. We analyze four different scenarios for each test case with clean and noisy (30 dB SNR) data for two different $N_{\text{train}}$ values. Instead of enforcing $\mathcal{L}_{\text{tc}}$ from the beginning, we enforce it after certain epoch $e_{s}$, a hyperparameter for training tcKAE. $\gamma_{\text{tc}}$. Up to $e_{s}$ we train without enforcing the consistency regularization, and switch to nonzero $\gamma_{\text{tc}}$ from $e_{s}$ onward. The weights for different loss components are one of the most crucial hyperparameters. Although, there are no fixed rule for selecting these weights, it is not entirely ad-hoc either. Since we are interested in forward prediction from the trained model, $\gamma_{\text{fwd}}$ and $\gamma_{\text{id}}$ are most crucial. In order to reduce the number of variables, we set $\gamma_{\text{id}}=1$ and vary other weights accordingly. Note that the advantage of tcKAE over DAE and cKAE comes at the cost of extra hyperparameters and increased training time. Training time, detailed training parameters, ablation study, and effect of $k_{\text{tm}}$ on prediction accuracy is provided in the supplementary material.
@@ Acknowledgements
I.N. and D.G. conceived the algorithm, I.N. developed the code and A.C. conducted the experiments, I.N. and A.C. analyzed the results. F.T., M.K., and D.G. advised I.N and A.C. and acquired the necessary funding. All authors reviewed the manuscript.

### v3

@@ Methods
The experiments were carried out on the Ohio Supercomputer’s (OSC) [37] Owens cluster, consisting of 842 Dell Nodes and 648 compute nodes with Dell PowerEdge C6320 two-socket servers, Intel Xeon E5-2680 v4 (Broadwell, 14 cores, 2.40 GHz) processors and 128 GB memory. As shown in Figure 4, we partition the dataset $\mathbf{X}$ into training ($\mathbf{X}_{\text{train}}$), validation ($\mathbf{X}_{\text{val}_{1}},\mathbf{X}_{\text{val}_{2}}$) and testing ($\mathbf{X}_{\text{test}}$) set. For a set of hyperparameters, the tcKAE architecture is trained to reduce the training loss $\mathcal{L}_{\text{tot}}$ as defined in 12. The set of hyperparameters which provides lowest validation loss is chosen to be the optimal set of hyperparameters for each case. The validation loss is computed as the average relative 2-norm error over $\mathbf{X}_{\text{val}_{2}}$, using predictions initialized from points in $\mathbf{X}_{\text{val}_{1}}$ as initial conditions. We analyze four different scenarios for each test case with clean and noisy (30 dB SNR) data for two different $N_{\text{train}}$ values. Instead of enforcing $\mathcal{L}_{\text{tc}}$ from the very first epoch of training, we enforce it after certain number of epochs $e_{s}$, which is a hyperparameter for training the tcKAE. That is, we keep $\gamma_{\text{tc}}=0$ for the first $e_{s}$ number of epochs. Up to $e_{s}$ we train without enforcing the consistency regularization, and switch to nonzero $\gamma_{\text{tc}}$ from $e_{s}$ onward. The look-ahead step $k_{tm}$ for enforcing temporal consistency is an important hyperparameter. Ideally, a larger $k_{tm}$ should lead to better accuracy, since we are looking further into the future. However, as seen in (11), increasing $k_{tm}$ also increases the number of terms in the summation for the temporal consistency loss. Since in practice we are using a numerical stochastic optimization routine, the increasing complexity (and possibly non-convexity) of the loss term increases the possibility of the optimization algorithm to get stuck in a local minima, thereby resulting in higher error and increased training time. As a result, in practice, we typically avoid selecting very large $k_{tm}$ unless there is a significant improvement in accuracy. The weights for different loss components are one of the most crucial hyperparameters. Although, there are no fixed rule for selecting these weights, it is not entirely ad-hoc either. Since we are interested in forward prediction from the trained model, $\gamma_{\text{fwd}}$ and $\gamma_{\text{id}}$ are most crucial. In order to reduce the number of variables, we set $\gamma_{\text{id}}=1$ and vary other weights accordingly. Note that the advantage of tcKAE over DAE and cKAE comes at the cost of extra hyperparameters and increased training time. Training time, detailed training parameters, ablation study, and effect of $k_{\text{tm}}$ on prediction accuracy is provided in the supplementary material (see Supplementary Tables 1-12).
Results
Baseline
Performance on Benchmark Datasets
In order to streamline the discussion, we adopt the following notation: the complete dataset is denoted by $\mathbf{X}\in\mathbb{R}^{N_{\text{dim}}\times N_{\text{tot}}}$, where $N_{\text{dim}}$ is the dimension of the state-space, and $N_{\text{tot}}$ is the total number of time samples  [^1] Note that we distinguish between time steps and time samples. By time step, we refer to the discrete time instances used by high-fidelity numerical solvers to generate data. This high-fidelity data is typically sampled at specific time steps and used for KAE training.. We denote the training set as $\mathbf{X}_{\text{train}}\in\mathbb{R}^{N_{\text{dim}}\times N_{\text{train}}}$, validation set as $\mathbf{X}_{\text{val}}\in\mathbb{R}^{N_{\text{dim}}\times N_{\text{val}}}$, and testing set as $\mathbf{X}_{\text{test}}\in\mathbb{R}^{N_{\text{dim}}\times N_{\text{test}}}$ with $N_{\text{train}}$, $N_{\text{val}}$ and $N_{\text{test}}$ being the number of time samples in training, validation and testing set respectively. Note that validation for long-term time-series prediction is nontrivial, and authors in [6] does not mention any validation set. However, for practical applications, validation is crucial, and we validate using the following method: We divide the validation set in two sets $\mathbf{X}_{\text{val}_{1}}\in\mathbb{R}^{N_{\text{dim}}\times N_{\text{val}_{1}}}$ and $\mathbf{X}_{\text{val}_{2}}\in\mathbb{R}^{N_{\text{dim}}\times N_{\text{val}_{2}}}$, separated by $\mathbf{X}_{\text{train}}$ as shown in Figure 4, where $N_{\text{val}}=N_{\text{val}_{1}}+N_{\text{val}_{2}}$. We chose $N_{\text{val}_{1}}=N_{\text{val}_{2}}=\frac{1}{4}N_{\text{train}}$ for all the datasets. The datapoints in $\mathbf{X}_{\text{val}_{1}}$ are used as initial conditions to predict over the span of $\mathbf{X}_{\text{val}_{2}}$. The error metric we use in this work is the 2-norm relative error defined (at $n^{\text{th}}$ time sample) by $\delta_{n}=\frac{||\hat{\mathbf{x}}_{n}-\mathbf{x}_{n}||_{2}}{||\mathbf{x}_{n}||_{2}}$, where $\mathbf{x}$ is the original state and $\hat{\mathbf{x}}$ is the predicted state. The validation loss is determined by calculating the average relative 2-norm error across $\mathbf{X}_{\text{val}_{2}}$, where predictions are generated using initial conditions from points in $\mathbf{X}_{\text{val}_{1}}$. The test error is calculated as the 2-norm relative error in the extrapolation region. To generate an error bar, we use the variance quantified by the 90% confidence interval, defined by the 5th and 95th percentiles of the error. This provides a robust summary of the typical error range by excluding the extreme 5% values at both ends of the distribution. We follow the statistical analysis approach used in [6]. Specifically, we take the first $N_{s}$ time steps ($N_{s}=N_{\text{train}}$) from the testing set and use the corresponding states as initial conditions. The trained model is then used to predict up to a fixed number of time steps. For each initial condition, we compute the 2-norm error by comparing the predictions with the ground truth. These errors are then used to generate the confidence interval. The reported results represent the average of neural network training outcomes across 10 different random seeds.
(a) $N_{\text{train}}=32$, Clean.
(b) $N_{\text{train}}=32$, 30 dB SNR.
(c) $N_{\text{train}}=64$, Clean.
(d) $N_{\text{train}}=64$, 30 dB SNR.
(a) $N_{\text{train}}=28$, Clean.
(b) $N_{\text{train}}=28$, 30 dB SNR.
(c) $N_{\text{train}}=52$, Clean.
(d) $N_{\text{train}}=52$, 30 dB SNR.
Undamped Pendulum
$N_{\text{train}}=32$ $N_{\text{train}}=64$
Method Clean 30 dB SNR Clean 30 dB SNR
DAE 34.411 (20.12) 35.055 (18.83) 13.884 (12.10) 12.421 (9.75)
cKAE 26.415 (18.52) 23.202 (15.71) 12.388 (11.82) 11.316 (9.75)
tcKAE 12.561 (10.20) 15.147 (10.92) 7.491 (6.36) 9.826 (7.91)
Oscillating electron beam
For the electron beam oscillation, one period approximately comprises of 25 time samples. In order to assess the performance of the Koopman models for limited dataset, we train the models for $N_{\text{train}}=28$ and $52$, each for clean and noisy (30 dB SNR) data. For $N_{\text{train}}=28$; $N_{\text{val}_{1}}=N_{\text{val}_{2}}=7$, and for $N_{\text{train}}=52$; $N_{\text{val}_{1}}=N_{\text{val}_{2}}=13$. We extrapolate up to 1000 time samples for evaluating the test error. Growth of the relative prediction error over 1000 time samples is depicted in Figure 6. The snapshots of the electric field distribution and corresponding spatial error profiles for long-term prediction is shown in Figure 7. Note that the spatial error is calculated as following: $\delta(x,y)=\frac{|\mathbf{x}(x,y)-\hat{\mathbf{x}}(x,y)|}{||\mathbf{x}||_{2}}$. As can be seen from Figure 6 and Table 2, the advantage of tcKAE is prominent for limited dataset i.e. $N_{\text{train}}=28$.
Table 2: Error (%) comparison for different $N_{\text{train}}$ and noise levels for oscillating electron beam data. The quantities inside parenthesis show the width of 90% confidence interval.
$N_{\text{train}}=28$ $N_{\text{train}}=52$
Method Clean 30 dB SNR Clean 30 dB SNR
DAE 16.570 (4.22) 27.006 (6.20) 3.956 (0.87) 4.782 (0.74)
cKAE 14.958 (2.15) 21.901 (8.28) 3.370 (0.61) 3.900 (0.64)
tcKAE 4.003 (0.73) 6.575 (1.23) 1.489 (0.46) 3.376 (0.52)
(a) $N_{\text{train}}=28$, Clean.
(b) $N_{\text{train}}=28$, 30 dB SNR.
(c) $N_{\text{train}}=40$, Clean.
(d) $N_{\text{train}}=40$, 30 dB SNR.
Flow past cylinder
Flow past cylinder is another common dynamical system used for benchmarking data-driven models. The simulation was carried out using MATLAB’s FEAtool (Figure 3b). A cylinder with diameter 0.1 m is located at a height of 0.2 m. The fluid is characterized by its density $\rho=1$ Kg/m3, and dynamic viscosity $\mu=0.001$ Kg/m s. The flow is unsteady with a maximum velocity of 1 m/s and mean velocity $\frac{2}{3}$ of the maximum velocity. The simulation is run for 80 s until the steady state is achieved, and the horizontal $u$ component of the velocity is probed at every 0.02 s starting from 20 s. The complete dataset $\mathbf{X}\in\mathbb{R}^{2647\times 3001}$ consists of 3001 time samples. More details regarding the simulation setup is presented in Methods: Simulation Setup section. Growth of relative error over 2500 time samples is shown in Figure 8. Predicted velocity field snapshots and corresponding error profile for long-term prediction is shown in Figure 9. Here also we observe that tcKAE provides significant advantage for limited training data.
Table 3: Error (%) comparison for different $N_{\text{train}}$ and noise levels for flow past cylinder data. The quantities inside parenthesis show the width of 90% confidence interval.
$N_{\text{train}}=28$ $N_{\text{train}}=40$
Method Clean 30 dB SNR Clean 30 dB SNR
DAE 10.182 (2.24) 20.724 (3.60) 1.695 (0.18) 10.775 (1.29)
cKAE 7.153 (1.06) 18.376 (5.58) 1.616 (0.17) 10.933 (1.24)
tcKAE 3.808 (0.45) 13.311 (0.53) 0.874 (0.11) 9.275 (0.94)
In case of flow past cylinder, since one cycle approximately contains 29 time samples, we consider $N_{\text{train}}=28$ and $40$. For $N_{\mathrm{train}}=28$, we take $N_{\text{val}_{1}}=N_{\text{val}_{2}}=7$, and for $N_{\mathrm{train}}=40$, $N_{\text{val}_{1}}=N_{\text{val}_{2}}=10$ is taken. We evaluate the test error for 2500 time samples in the extrapolation region. Growth of relative error over 2500 time samples is shown in Figure 8. Predicted velocity field snapshots and corresponding error profile for long-term prediction is shown in Figure 9. Table 3 provides the relative (percentage) time-averaged test error. Here also we observe that tcKAE provides significant advantage for limited training data.
Discussion
The proposed tcKAE algorithm advances the state-of-the-art by enforcing multi-step temporal consistency in the latent space. Unlike prior KAE architectures such as consistent Koopman autoencoders (cKAE) [6], which rely on backward dynamics assumptions, tcKAE sidesteps this limitation by introducing a consistency regularization term in the forward direction. The ablation study in the supplementary material further demonstrates that, even without the backward and consistency losses, the combination of forward, identity, and temporal consistency losses achieves comparable or even superior accuracy to cKAE. This innovation strengthens the Koopman invariance of the latent space, enhancing robustness to limited date and noise, improving generalizability. Furthermore, empirical results demonstrate that tcKAE outperforms both cKAE and dynamic autoencoders (DAEs) [7] in terms of long-term prediction accuracy and stability, particularly for scenarios where we have limited training data. The proposed tcKAE not only achieved lower error margins in benchmark problems such as pendulum oscillation, oscillating electron beam, and flow past a cylinder, but also demonstrated narrower confidence intervals, thereby inspiring greater confidence in the predictions. It is important to note that the proposed tcKAE does not guarantee unconditional stability. The tcKAE predictions are more stable for long-term predictions compared to DAE and cKAE. However, the secular growth in relative error indicates that if we were to predict 10 or 20 times further, the large error would likely render the results useless for all the models.
Data Availability
The code for generating the data and results of this paper will be made available upon request. Any such request should be addressed to A.C. It is also partially available at this dynamically updated GitHub repository.
@@ Acknowledgements
I.N. and D.G. conceived the algorithm. I.N. developed the code. A.C. conducted the experiments. I.N. and A.C. analyzed the results. F.T., M.K., and D.G. advised I.N and A.C. and acquired the funding and resources. All authors reviewed the manuscript.